the weight of the object measured in the mars in 10 n where the observed gravitational acceleration is 3.40 m/ s^2 . determine the weight of the object in earth where the gravitational acceleration is 9. 81 m/s^2? solution


1
Expert's answer
2021-03-08T08:26:43-0500

Let's first find the mass of the object:


"m=\dfrac{W}{g}=\dfrac{10\ N}{3.4\ \dfrac{m}{s^2}}=2.94\ kg."

Finally, we can find the weight of the object on Earth:


"W=mg=2.94\ kg\cdot9.81\ \dfrac{m}{s^2}=28.84\ N."